# Heisenberg’s S-matrix program and Feynman’s divergence problem

### v1

@@ Abstract
Keywords: Generalized wave operators, deviation factors, divergence problems, power series, multiplicative integrals.
@@ 3 Secondary generalized scattering
1. It is well known that Coulomb type potentials generate infrared divergences (infrared catastrophe) . The problem has a long history starting in 1940 with the fundamental works of R. Feynman. Here, we construct a rigorous solution of the corresponding problem for a broad class of cases.
Remark 3.1
Interesting papers on the infrared case [12, 14] are closely related to the Faddeev-Kulish construction [15]. However, our approach is essentially different ((even without the secondary generalization modification considered here)). In greater detail, our approach and Faddeev-Kulish construction are compared in [26].
Recall the operator function
@@
Clearly, $S^{R}(t,\tau,\varepsilon)=S(t,\tau,\varepsilon)$ and $U(t,\varepsilon)=u(t)$ in the case $t,\tau\in[-1,1]$. We also have
$S^{R}(0,0,{\varepsilon})=I.$ (3.16)
The next theorem easily follows from the relations of this section.
@@
$S^{R}(t,\tau,\varepsilon)=S^{R}(t,1,\varepsilon)S^{R}(1,0,\varepsilon)S^{R}(0,-1,\varepsilon)S^{R}(-1,\tau,\varepsilon).$ (3.17)
@@
$S^{R}(t,0,\varepsilon)=\overset{t}{\overset{\curvearrowleft}{\underset{0}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}}=S^{R}(t,1,\varepsilon)S^{R}(1,0,\varepsilon)\qquad(t\geq 1).$ (3.18)
In a similar way, we derive from (3.14) the equality
$S^{R}(0,\tau,\varepsilon)=\overset{0}{\overset{\curvearrowright}{\underset{\tau}{\int}}}e^{i\varepsilon{U(s,\varepsilon)ds}}=S^{R}(0,-1,\varepsilon)S^{R}(-1,\tau,\varepsilon)\qquad(\tau\leq-1).$ (3.19)
Finally, relations (3.13) and (3.14) yield for $t\geq 1$ and $\tau\leq-1$ that
$S^{R}(t,\tau,\varepsilon)=\overset{t}{\overset{\curvearrowleft}{\underset{0}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}}\overset{0}{\overset{\curvearrowright}{\underset{\tau}{\int}}}e^{i\varepsilon{U(s,\varepsilon)ds}}=S^{R}(t,0,\varepsilon)S^{R}(0,\tau,\varepsilon).$ (3.20)
Formula (3.17) immediately follows from (3.18)–(3.20). $\Box$
Formula (3.1) implies that the operators $S(t,\tau,{\varepsilon})$ are unitary for the case, where the self-adjoint operators $A_{0}$ and $A_{1}$ exist. In our general case, it follows from (3.6) and (3.10) that $W_{0}(t)$ is unitary. Hence, (3.15) implies that $U(t,{\varepsilon})$ is self-adjoint. Now, relations (3.17)–(3.19) yield the unitarity assertion:
@@
$S^{R}(t,\tau,\varepsilon),\,\,S^{R}(t,1,\varepsilon),\,\,S^{R}(1,0,\varepsilon),\,\,S^{R}(0,-1,\varepsilon),\,\,S^{R}(-1,\tau,\varepsilon),$ (3.21)
@@
In fact, it is easy to see that all the operators $S^{R}(t,\tau)$ are unitary (without the restriction $t\geq 1$ and $\tau\leq-1$). However, only the unitarity of the operators given in (3.21) is used further in the text.
Next we formulate the main result of this section, which follows from Theorem 3.4, Propositions 3.5 and B.1, and Corollary B.2.
@@
$S^{R}(+\infty,-\infty,\varepsilon)=S^{R}(+\infty,1,\varepsilon)S(1,0,\varepsilon)S(0,-1,\varepsilon)S^{R}(-1,-\infty,\varepsilon),$ (3.22)
@@
$S^{R}(+\infty,1,\varepsilon)=\overset{+\infty}{\overset{\curvearrowleft}{\underset{1}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}},\quad S^{R}(-1,-\infty,\varepsilon)=\overset{-1}{\overset{\curvearrowright}{\underset{-\infty}{\int}}}e^{i\varepsilon{U(s,\varepsilon)ds}}.$ (3.23)
@@
It is proved (see [2] and the review [6] from [2] in particular) that in many cases the Feynman amplitudes have the polylogarithmic structure as in (3.5) and (3.32).
@@ Appendix B Multiplicative integrals
then the corresponding right multiplicative integrals converge by norm for $a\to-\infty:$
$\overset{b}{\overset{\curvearrowright}{\underset{a}{\int}}}e^{F(t)dt}{\Longrightarrow}\overset{b}{\overset{\curvearrowright}{\underset{-\infty}{\int}}}e^{F(t)dt}\qquad(a\to-\infty).$ (B.7)

### v2

@@ Abstract
Keywords: Generalized wave operator, deviation factor, secondary generalized scattering operator, infrared divergence, ultraviolet divergence.
@@ 3 Secondary generalized scattering
1. It is well known that Coulomb type potentials generate infrared divergences (infrared catastrophe) . The problem has a long history starting in 1940 with the fundamental works of R. Feynman. Here, we construct a rigorous solution of the corresponding problem for a broad class of cases. Recall the operator function
@@
Clearly, we have
$S^{R}(0,0,{\varepsilon})=I\quad{\mathrm{and}}\quad S^{R}(t,\tau,\varepsilon)=S(t,\tau,\varepsilon),\quad U(t,\varepsilon)=u(t)$ (3.16)
for $t,\tau\in[-1,1]$. The next theorem easily follows from the relations of this section.
@@
$S^{R}(t,\tau,\varepsilon)=S^{R}(t,1,\varepsilon)S^{R}(1,-1,\varepsilon)S^{R}(-1,\tau,\varepsilon).$ (3.17)
@@
$S^{R}(t,\tau,\varepsilon)=\overset{t}{\overset{\curvearrowleft}{\underset{\tau}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}}\qquad(t\geq\tau).$ (3.18)
Hence, for $t\geq 1$ and $\tau\leq-1$ we have
$S^{R}(t,\tau,\varepsilon)=\overset{t}{\overset{\curvearrowleft}{\underset{1}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}}\overset{1}{\overset{\curvearrowleft}{\underset{-1}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}}\overset{-1}{\overset{\curvearrowleft}{\underset{\tau}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}}.$ (3.19)
Formula (3.17) immediately follows from (3.18) and (3.19). $\Box$
Formula (3.1) implies that the operators $S(t,\tau,{\varepsilon})$ are unitary for the case, where the self-adjoint operators $A_{0}$ and $A_{1}$ exist. In our general case, it follows from (3.6) and (3.10) that $W_{0}(t)$ is unitary. Hence, (3.15) implies that $U(t,{\varepsilon})$ is self-adjoint. Now, relation (3.18) yields the unitarity assertion:
@@
$S^{R}(t,\tau,\varepsilon),\,\,S^{R}(t,1,\varepsilon),\,\,S^{R}(1,-1,\varepsilon),\,\,S^{R}(-1,\tau,\varepsilon),$ (3.20)
@@
In fact, it is easy to see that all the operators $S^{R}(t,\tau,{\varepsilon})$ are unitary (without the restriction $t\geq 1$ and $\tau\leq-1$). However, only the unitarity of the operators given in (3.20) is used further in the text.
Remark 3.5
Clearly, formula (3.17) ((under conditions of Theorem 3.3)) may be rewritten as
$S^{R}(t,\tau,\varepsilon)=S^{R}(t,1,\varepsilon)S^{R}(1,0,\varepsilon)S^{R}(0,-1,\varepsilon)S^{R}(-1,\tau,\varepsilon),$ (3.21)
which is sometimes more convenient in the case of a special behaviour at zero. We also have
$S^{R}(t,0,\varepsilon)=S^{R}(t,1,\varepsilon)S^{R}(1,0,\varepsilon).$ (3.22)
All the factors in (3.21) and (3.22) are again unitary.
Next we formulate the main result of this section, which follows from Theorem 3.3, formula (3.16), Propositions 3.4 and B.1, and Corollary B.2.
@@
$S^{R}(+\infty,-\infty,\varepsilon)=S^{R}(+\infty,1,\varepsilon)S(1,-1,\varepsilon)S^{R}(-1,-\infty,\varepsilon),$ (3.23)
@@
$S^{R}(+\infty,1,\varepsilon)=\overset{+\infty}{\overset{\curvearrowleft}{\underset{1}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}},\quad S^{R}(-1,-\infty,\varepsilon)=\overset{-1}{\overset{\curvearrowleft}{\underset{-\infty}{\int}}}e^{-i\varepsilon{U(s,\varepsilon)ds}}.$ (3.24)
@@ 4 Ultraviolet divergence
Interesting results on the asymptotic behaviour of the Feynman integrals (4.2) are derived in [6] (see also some other papers and references in [2]).
@@ 2. The commutation relation
Generalized wave operators were also used in [15]. Interesting papers on the infrared case [12, 14] are closely related to the Faddeev-Kulish construction [15].
@@ Appendix B Multiplicative integrals
then the corresponding left and right multiplicative integrals converge by norm for $a\to-\infty:$
$\overset{b}{\overset{\curvearrowleft}{\underset{a}{\int}}}e^{F(t)dt}{\Longrightarrow}\overset{b}{\overset{\curvearrowleft}{\underset{-\infty}{\int}}}e^{F(t)dt},\qquad\overset{b}{\overset{\curvearrowright}{\underset{a}{\int}}}e^{F(t)dt}{\Longrightarrow}\overset{b}{\overset{\curvearrowright}{\underset{-\infty}{\int}}}e^{F(t)dt}\qquad(a\to-\infty).$ (B.7)